# Structured Packing in LLM Training Improves Long Context Utilization

### v6

@@ Abstract
Recent developments in long-context large language models have attracted considerable attention. Yet, their real-world applications are often hindered by ineffective context information use. This work shows that structuring training data to increase semantic interdependence is an effective strategy for optimizing context utilization. To this end, we introduce Structured Packing for Long Context (SPLiCe), a method for creating training examples by using information retrieval methods to collate mutually relevant documents into a single training context. We empirically validate SPLiCe on large $3$B and $7$B models, showing perplexity improvements and better long-context utilization on downstream tasks. Remarkably, already relatively short fine-tuning with SPLiCe is enough to attain these benefits. Additionally, the comprehensive study of SPLiCe reveals intriguing transfer effects such as training on code data leading to perplexity improvements on text data.
Machine Learning, Language Models, ICML
@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023b; Bai et al., 2023) have transformed the field of AI. Recently, the field has observed the rise of Long-Context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic, 2023; OpenAI, 2023a). However, their ability to process long context is not always as effective as one hopes for. Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023a). Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al., 2023), even on simple synthetic retrieval tasks (Li et al., 2023a). Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components, such as the architecture and training objectives, unchanged. The broad question is how to organize training data to enhance long context capabilities? Such perspective has received some attention recently (Levine et al., 2022; Chan et al., 2022; Shi et al., 2023b), but the problem is far from being solved. The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization. We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training examples by using retrieval (e.g., BM25 or repository structure) to collate mutually relevant documents into a single training context.
@@
We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2 and $7$Bv2 (Geng & Liu, 2023) for only $2$B–$6$B tokens already brings perplexity reduction. This reduction translates to substantial improvements in handling long-context information in downstream tasks that require retrieval and in-context learning. These tasks include TREC (Li & Roth, 2002; Hovy et al., 2001), DBpedia (Lehmann et al., 2015), Qasper (Shaham et al., 2022; Dasigi et al., 2021), HotPotQA (Yang et al., 2018), and the lost-in-the-middle benchmark (Liu et al., 2023). We perform a comprehensive study of the design choices and properties of SPLiCe, showing that the acquired long-context capabilities transfer between modalities, such as code and text.
@@
  • We fine-tune OpenLLaMA $3$Bv2 and OpenLLaMA $7$Bv2 (Geng & Liu, 2023) using SPLiCe, showing that it improves perplexity on long context evaluation, in-context learning ability, and ability to retrieve relevant data from the context.
@@ 2 Method
SPLiCe is a method for constructing training examples that enhances the LLMs long-context utilization. This translates to improved performance on a range of tasks like in-context learning, question answering, in-context information retrieval, and long context language modeling. We make the source code available to ensure the reproducibility of our results. (For this anonymous submission we share the zip file with our code.)
@@
  Input:    $D$: document corpus   $k$: breadth hyper-parameter   $L$: maximum length of returned training example   RETRIEVE: retrieval method to use, e.g., BM25   ORDER: ordering method, e.g., id, or random shuffle
@@
           {RETRIEVE uses a precomputed index, and may return documents that are already in $C$}
@@
Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al., 2022). However, it is an open question how to achieve such benefits in pre-training or fine-tuning. The primary difficulty comes from the fact that long-range dependencies are rare in training data (de Vries, 2023) and diminish with distance. It is thus unlikely that a model will learn to utilize long context without more explicit guidance.
Recent studies indicate that structuring data, i.e., going beyond the i.i.d. paradigm, might be beneficial or even necessary to achieve good long-context performance. (Levine et al., 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training example. (Chan et al., 2022) indicates that specific distributional properties of the training data have a significant impact on the model’s in-context performance. Finally, (Shi et al., 2023b) shows that training on structured data improves the performance of LLMs on long-context tasks.
@@
We propose SPLiCe, a method that creates training examples by building a tree of documents, see Algorithm 1 and Figure 1 for a general overview.
SPLiCe starts by picking a random document from the dataset to create a single-node tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus. The final sequence consists of the tree flattened according to a certain tree traversal strategy.
The hyperparameter $k$ introduces flexibility and allows to interpolate between modes of retrieval. Indeed, for $k=1$ SPLiCe creates a path of related examples simulating a long document. For larger values of $k$, SPLiCe produces examples that are similar to the ones used by retrieval augmented models, e.g., (Borgeaud et al., 2022). Note that the dataset must be reasonably curated for SPLiCe to work efficiently. For example, a low duplication rate incentivizes the model to utilize long context beyond simple copying.
@@
  • SPLiCeBM25: based on BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
  • SPLiCeCont: based on Contriever-MSMARCO (Cont) (Izacard et al., 2022), a recently proposed retrieval method that uses a transformer model to rank documents based on their similarity to a query.
SPLiCe computational efficiency Given the dataset sizes used in training LLMs, the computational efficiency plays a non-trivial role. SPLiCeRepo is the fastest and easy to implement but requires additional directory structure, i.e., is not applicable to general web data. SPLiCeBM25 requires quadratic time index and needs dataset chunking to be applicable to large datasets, see Section 3.3. SPLiCeCont requires calculating embeddings for each document and retreval based on the cosine similarity. The latter step can be done efficiently using fast approximate inner-product search, e.g., Faiss (Johnson et al., 2017).
Baseline The standard approach, commonly used in LLM training pipelines, is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context. This is known as example packing (Brown et al., 2020).
Burstiness SPLiCe conjecturally falls into the framework presented in (Chan et al., 2022). This paper shows that the distributional properties of the training data affect the in-context capabilities of transformer models. In particular, it indicates the importance of ”burstiness”, i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence . In Appendix K, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) compared to the baseline.
3 Experiments with Medium-Scale Models
In this section, we conduct a comprehensive examination of the impact of document packing on long-context performance, using medium-scale models (see Section 4 for large models results). We demonstrate that structuring data with SPLiCe is a generally applicable approach to improve long context performance. We study the performance in Section 3.2 and design choices of SPLiCe in Section 3.3.
@@ 3.1 Experimental Setup
In this part, we employ $270$M models and the following protocol. Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer, 2023). Subsequently, we continue training using $1$B tokens with an extended context length of $32$K on a mixture of the original RedPajama data (TogetherComputer, 2023) and long context data created using SPLiCe/Baseline. We employ the Focused Transformer (FoT) objective (Tworkowski et al., 2023) for context extension (unless stated otherwise). This approach is motivated by practical factors, viz. training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023). Loosely inspired by (Ouyang et al., 2022; Rozière et al., 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture. This aims to prevent the model from overfitting to artificially created long documents.
In the evaluation, we measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets employing a context length of $32$K. The selection of these datasets is motivated by their inclusion of documents that can benefit from long-context information as demonstrated in (Chen et al., 2023) and (Li et al., 2023b). For example, functions are often re-utilized, and similar terms are employed across papers. We exclude documents with fewer than $32$K tokens and truncate those exceeding this length.
We refer to Appendix A for a detailed description of the model architecture. For information regarding the training and evaluation data, see Appendix G.
@@ 3.2 Experimental Results
In this section, we compare SPLiCe against Baseline. We test two retrieval approaches for SPLiCe: BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), and Contriever-MSMARCO (Izacard et al., 2022)) denoted as SPLiCeBM25 and SPLiCeCont respectively. In SPLiCeBM25 retrieval takes into account the whole content of the document, whereas in SPLiCeCont we use the first $512$ tokens and Faiss (Johnson et al., 2017) for fast approximate inner-product search (see Appendix H for details). By SPLiCeRepo, we denote the approach using the directory structure of code repositories. We also test our method with $64$K context length and report the results in Table 19 in Appendix J. In those experiments, we use $k=1$; see Section 3.3 for more details.
The results are presented in Table 3, Table 3, and Table 3, from which we draw the following conclusions.
Structuring data improves performance. We observe (see Table 3) that all variants of SPLiCe outperform Baseline, often significantly, regardless of the training and eval dataset. This indicates that the structure of the data is important for long-context training.
The positive effects transfer between domains. We observe (see Table 3) positive effects when training on code and evaluating on arXiv and when training on web data and evaluating on code. This indicates that the positive effects of SPLiCe transfer between domains. We conjecture that better data mixtures could bring further benefits.
Retrieval method is of secondary importance. We observe (see Table 3) that the choice of retrieval method has a minor impact on the results. Typically, SPLiCeBM25 is slightly better than SPLiCeCont. On the code data, SPLiCeRepo is on par with SPLiCeBM25. We recall that SPLiCeRepo can be applied only to code data, whereas SPLiCeBM25 and SPLiCeCont are more general. In practice, we often observe that SPLiCeCont is faster than SPLiCeBM25 due to the use of Faiss.
@@
Long Context Method arXiv Code Code &

### v7

@@ Abstract
Recent advancements in long-context large language models have attracted significant attention, yet their practical applications often suffer from suboptimal context utilization. This study investigates structuring training data to enhance semantic interdependence, demonstrating that this approach effectively improves context utilization. To this end, we introduce the Structured Packing for Long Context (SPLiCe) method, which utilizes retrieval to collate mutually relevant documents into long and coherent training examples. We validate SPLiCe empirically across models of varying sizes — $3$B, $7$B, and $13$B — achieving improved performance in long-context tasks, such as Qasper and HotpotQA. Remarkably, even brief fine-tuning with SPLiCe is sufficient to realize these benefits. Additionally, SPLiCe effectively mitigates the ‘lost-in-middle’ phenomenon often observed in large models. Our comprehensive analysis of SPLiCe explores its design choices and reveals intriguing transfer effects; for instance, training on programming code enhances performance on natural language tasks. 11footnotetext: University of Warsaw, 2Ideas NCBR, 3University of Edinburgh, 4Google DeepMind, 5Institute of Mathematics, Polish Academy of Sciences, 6deepsense.ai
@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023b; Bai et al., 2023) have transformed the field of AI. Recently, the field has observed the rise of long context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic, 2023; OpenAI, 2023a; Gemini Team, 2024). However, their ability to process long context is not always as effective as one hopes for. Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023a). Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al., 2023), even on simple synthetic retrieval tasks (Li et al., 2023a). Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components, such as the architecture and training objectives, unchanged. The broad question is how to organize training data to enhance long context capabilities? Such perspective has received some attention recently (Levine et al., 2022; Chan et al., 2022; Shi et al., 2023b), but the problem is far from being solved. The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization. We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training examples by using retrieval (e.g., BM25, Contriver or repository structure) to collate mutually relevant documents into a single training context.
@@
We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2, $7$Bv2 (Geng and Liu, 2023) and CodeLlama $13$B (Rozière et al., 2023) with mere $2$B–$6$B tokens already brings substantial improvements in handling long context information in downstream tasks that require retrieval and in-context learning. These tasks include Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022), lost-in-the-middle benchmark (Liu et al., 2023), HotPotQA (Yang et al., 2018), TREC (Li and Roth, 2002; Hovy et al., 2001), and DBpedia (Lehmann et al., 2015). We perform a comprehensive study of the design choices and properties of SPLiCe, showing, in particular, that the acquired long context capabilities transfer between modalities, such as code and text. SPLiCe also helps to retain and in some cases even improve performance on short context benchmarks like GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and TREC (Li and Roth, 2002; Hovy et al., 2001).
@@
  • We fine-tune OpenLLaMA $3$Bv2, OpenLLaMA $7$Bv2 (Geng and Liu, 2023) and CodeLlama $13$B (Rozière et al., 2023) using SPLiCe, showing that it improves downstream performance on long context tasks.
@@ 2 Method
SPLiCe is a method for constructing training examples that enhance the LLMs’ long context utilization. This translates to improved performance on a range of tasks like in-context learning, question answering, in-context information retrieval, and long-context language modeling.
@@
Input:
   $D$: document corpus
   $k$: breadth hyper-parameter
   $L$: maximum length of returned training example
   RETRIEVE: retrieval method to use, e.g., BM25
   ORDER: ordering method, e.g., identity, or random shuffle
@@
         {RETRIEVE uses a precomputed index and may return documents that are already in $C$}
@@
Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al., 2022). However, it is an open question how to achieve such benefits in pre-training or fine-tuning. The primary difficulty comes from long-range dependencies being rare in training data (de Vries, 2023) and diminishing with distance. Thus, it is unlikely that a model will learn to utilize long context without more explicit guidance.
Recent studies indicate that structuring data, i.e., going beyond the i.i.d. paradigm, might be beneficial or even necessary to achieve good long-context performance. (Levine et al., 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training example. Chan et al. (2022) indicates that specific distributional properties of the training data have a significant impact on the model’s in-context performance. Finally, concurrently to our work (Shi et al., 2023b) shows that pre-training on structured data improves the performance of LLMs.
@@
SPLiCe starts by picking a random document from the dataset to create a root of a tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus. The final sequence consists of the tree flattened according to a certain tree traversal strategy; see Algorithm 1.
The hyperparameter $k$ introduces flexibility, enabling interpolation between different modes of retrieval. Specifically, when $k=1$, SPLiCe simulates a long document by creating a path of related examples. For larger $k$ values, SPLiCe generates examples akin to those used in retrieval-augmented models, e.g., (Borgeaud et al., 2022). It is essential for the dataset to be well-curated to ensure the efficient functioning of SPLiCe. For instance, maintaining a low duplication rate encourages the model to leverage long context beyond merely copying. In this work, we employ the StarCoder (Li et al., 2023b) dataset, which has been deduplicated using the pipeline described in (Allal et al., 2023).
@@
  • SPLiCe BM25 : based on BM25 (Robertson and Zaragoza, 2009; Bassani, 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
  • SPLiCe Cont : based on Contriever-MSMARCO (Cont) (Izacard et al., 2022), a recently proposed retrieval method that uses a transformer to rank documents based on their similarity to a query.
Baselines One standard approach, commonly used in LLM training pipelines, is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context. This is known as example packing (Brown et al., 2020). Another popular approach restricts sampled documents to the same meta-class (for example, Wikipedia article, source code) (Groeneveld et al., 2024; Zhao et al., 2024). We compare SPLiCe against both.
SPLiCe computational efficiency Given the dataset sizes used in training LLMs, computational efficiency plays a non-trivial role. SPLiCe Repo is the fastest and easiest to implement but requires additional directory structure, i.e., it does not apply to general web data. SPLiCe BM25 uses a bag of words BM25 method that lacks deeper semantic encoding. SPLiCe Cont requires calculating embeddings for each document and retrieval based on the vector inner-product. The latter step can be done efficiently using fast approximate inner-product search, e.g., Faiss (Johnson et al., 2017).
Burstiness SPLiCe conjecturally falls into the framework presented in Chan et al. (2022), which shows that the distributional properties of the training data affect the in-context capabilities of transformer models. In particular, it indicates the importance of "burstiness", i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence. In Appendix K, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) compared to the example packing baseline.
3 Experiments
In this section, we show that SPLiCe improves the long context performance of large-scale language models. To this end, we use $3$B, $7$B, and $13$B parameter models and evaluate on tasks that test in-context learning (Section 3.2.1), question answering (Section 3.2.2), and in-context information retrieval capabilities (Section 3.2.3). Additionally, we provide perplexity results in Appendix E. Next, we show that SPLiCe can improve the core model capabilities by testing the short context performance, see Section 3.2.5. Finally, in Section 3.3, we tune over 30 medium ($270$M parameters) models using different data mixtures and SPLiCe parameters, to analyze its design choices.
An important finding of our work is that presented improvements occur during a relatively short, compared to pre-training, fine-tuning. To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B and $13$B models were tuned on $2$B tokens. The $3$B and $7$B start from OpenLLaMAv2 (Geng and Liu, 2023) that was pre-trained on $1$T tokens, the $13$B parameter models start from CodeLlama (Rozière et al., 2023) that was pre-trained on $2.5$T tokens.
@@ 3.1 Experimental setup
For $3$B experiments, we continue training on a 50/50 mixture of RedPajama, prepared in the standard way, and C prepared using SPLiCe BM25 . For $7$B and $13$B ones, we continue the training on a 50/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCe BM25 . StackExchange is part of RedPajama TogetherComputer (2023), and C data come from StarCoder (Li et al., 2023b). Including the standard RedPajama aims to prevent the model from overfitting to artificially created long documents. We refer to Appendix A for a detailed description of the model architectures.
We train with $32$K context length. We use a batch size of $256$K tokens per step and the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng and Liu, 2023). For $13$B model we use batch of $512$K tokens per step.
Regarding the context extensions method, we use FoT (Tworkowski et al., 2023), and the CodeLlama (CL) (Rozière et al., 2023) approach. More explicitly, we test eight models:
[…]
where Baseline denotes the standard (example packing) data preparation method where context is created by sampling random documents from the corpus and separating them with BOS/EOS tokens. We provide results regarding the within-domain sampling baseline in Appendix C.2. If not stated otherwise, in SPLiCe we use $k=1$ and the identity permutation as Order in the Algorithm 1. Hyperparameter details can be found in Appendix A.
@@ 3.2 Experimental results
3.2.1 In-Context learning
To evaluate the improvements of in-context learning abilities, we use TREC (Li and Roth, 2002; Hovy et al., 2001) and DBpedia (Lehmann et al., 2015), which are text classification tasks. For TREC, to test the long context utilization, we vary the number of in-context examples in $\{780,1560\}$ (which correspond roughly to $\{16,32\}$K context length). We sample these examples multiple times and measure classification accuracy on the test set. In Table 1, we observe that SPLiCe improves the average performance across different context lengths and model sizes. We follow a very similar protocol for DBpedia ($\{16,32\}$K context length corresponding to $\{190,380\}$ in-context examples), confirming that SPLiCe improves the context utilization. In Appendix I, we study the distribution of improvements with respect to the choice of the in-context examples, showing the stochastic domination of SPLiCe over the baseline.
Table 1: We test the average classification performance on TREC (Li and Roth, 2002; Hovy et al., 2001). We average across $50$ sets of in-context examples for $3$B models, $10$ sets for $7$B models, and $5$ sets for $13$B models. We use $\Delta$ [confidence interval] to denote the mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix I). For DBpedia, we average results across $40$ sets of in-context examples for $3$B and $7$B models and $5$ for $13$B ones. Due to the size of the DBpedia evaluation dataset, for each set of in-context examples, we subsample a subset of $500$ elements of the evaluation set.
Task TREC DBpedia
Model Context Baseline SPLiCe $\Delta$ [conf interv] Baseline SPLiCe $\Delta$ [conf interv]
@@ 3.2.1 In-Context learning
$16$K 68.9 76.9 8.0 [6.9,9.3]{6.9,9.3]}6.9 , 9.3 ] 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4]}2.5 , 3.4 ]
[…]
$16$K 74.0 79.0 5.0 [3.4,6.0]{3.4,6.0]}3.4 , 6.0 ] 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5]}1.5 , 2.5 ]
[…]
$16$K 81.4 82.5 1.1 [0.2,1.6]{0.2,1.6]}0.2 , 1.6 ] 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3]}0.0 , 0.3 ]
[…]
$16$K 88.2 91.2 3.0 [1.9,3.5]{1.9,3.5]}1.9 , 3.5 ] 95.8 96.8 1.0 [0.6,1.5]{0.6,1.5]}0.6 , 1.5 ]
Table 2: We measure question answering over long input using Qasper (Dasigi et al., 2021) (2-shot setting) and HotPotQA (Yang et al., 2018) (10-shot setting). For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al., 2021). For HotPotQA, we average results across $7$ sets of in-context examples for $3$B and $7$B models and $2$ for $13$B ones, with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals. Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
Task QASPER HotPotQA
Model Context Baseline SPLiCe Context Baseline SPLiCe $\Delta$ [conf interv]
[…]
3.2.2 Question answering
We also show that SPLiCe models have improved question answering capabilities. To this end, we use Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022) and HotPotQA (Yang et al., 2018). Qasper contains questions regarding research papers provided as a part of the input context. In Table 2, we observe that SPLiCe improves the two-shot performance. Similarly, we observe improvements on HotPotQA, which tests the ability to aggregate knowledge across multiple documents. We stress that those results measure few-shot performance, as neither of the datasets was used during training.
3.2.3 Lost in the middle
The key retrieval task introduced in (Liu et al., 2023) has become a routine diagnostic tool to study context capabilities. Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key. Even though very simple, the task proved to be challenging for many open-source language models. Specifically, LMs tend to achieve better performance when the relevant information is either in the beginning or the end of the input context, while struggling in the middle. See Appendix D for the task details.
In Figure 2, we present the key retrieval performance of $7$B models trained with CodeLlama (Rozière et al., 2023) context extension method. We note that despite relatively short tuning, SPLiCe significantly helps on hard-to-retrieve positions (the position of the key in the range $[25,100]$). We present more details in Appendix D.
Figure 2: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens). The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions in the middle. The details about this task can be found in Appendix D. Each position was evaluated using $500$ examples.
3.2.4 Ordering of examples
We typically use the identity ordering as Order in Algorithm 1 to merge documents into a single context, as we found that in most cases, it performs best. In Appendix F, we found that random shuffling is slightly better in some cases. Specifically, this is the case of large $7$B CodeLlama models, see Table Table 18. We hypothesize that random ordering forces the model to make use of the full space of the RoPe positional encoding. Whereas the identity ordering, also used in (Shi et al., 2023b), skews the model toward paying more attention to fragments of text that are not too far away. This suggests an interesting research direction on the intersection of data preparation and positional embeddings.
3.2.5 Short context evaluation
An undesirable side effect of fine-tuning for extended context (e.g., with SPLiCe) could be performance degradation on shorter contexts. To test this, we evaluate the short context performance of our models using GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and TREC (Li and Roth, 2002; Hovy et al., 2001). We present the results in Table 3. We note that for smaller models, both SPLiCe and Baseline show some short context degradation, whereas the larger $13$B CodeLlama Model SPLiCe brings significant improvements in short context performance on GSM8K ($+1.7$), TREC ($+3.3$) and maintains the MMLU performance. We provide an additional evaluation on code in Appendix L.
Table 3: We evaluate our models on MMLU ($5$-shot), GSM8K ($8$-shot) and TREC ($90$-shot). We provide an additional comparison with their starting checkpoint. For $7$B case, we additionally compare with a model tuned with $2$k context length on the same data. For each task, we highlight best results up to $1$ point.
Model MMLU GSM8K TREC
STEM Humanities Social Other Mean
$3$B Starting Chkpt 24.4 25.6 26.3 27.8 25.9 4.3 58.4
[…]
$7$B Starting Chkpt 33.6 42.1 46.0 44.5 40.8 8.0 59.6
$7$B 2K Tuned 33.7 40.8 44.6 41.1 39.4 8.4 62.3
[…]
$13$B Starting Chkpt 36.6 48.2 50.6 45.4 44.3 21.4 78.7
[…]
3.3 Detailed study with medium models
In Table 5 and Table 5, we present a comprehensive examination of the impact of document packing on long-context performance using $270$M parameter models. Using the smaller size allows us to train over $30$ models and, thus, more accurately estimate the impact of various design choices.
Training protocol
Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer, 2023). Subsequently, we continue training using $1$B tokens with the context extended to $32$K on a mixture of the original RedPajama data (TogetherComputer, 2023) and long context data created using SPLiCe/Baseline. The amount of pre-training tokens is based on scaling laws from (Hoffmann et al., 2022) and constants calculated for GPT-like models in (Karpathy, 2022).
We employ the Focused Transformer (FoT) objective Tworkowski et al. (2023) for context extension (unless stated otherwise). This approach is motivated by practical factors, viz. training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023). Loosely inspired by (Ouyang et al., 2022; Rozière et al., 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture. We also check how SPLiCe works with other context extension methods in Appendix B.2, and 64K context length in Appendix J.
Evaluation
We measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets employing a context length of $32$K. The selection of these datasets is motivated by the fact that they can benefit from long-context information as demonstrated in (Chen et al., 2023; Li et al., 2023b). For example, functions are often re-utilized, and similar terms are employed across papers. We exclude documents with fewer than $32$K tokens and truncate those exceeding this length. In Appendix B.1 we evaluate our models using short context data confirming no performance degradation with respect to the base model. For information regarding the training and evaluation data, see Appendix G.
[…]
Altered Method arXiv Code Code &